Find all the real roots of equation x^4 + 2x^2 − 24 = 0.

Answer by ewatrrr(24785)


Hi
x^4 + 2x^2 − 24 = 0.
Find all the real roots of equation
completing the square
(x^4 + 2x^2 + 1) - 1 − 24 = 0.
(x^2 +1)^2 -25 = 0
(x^2 +1)^2 = 25
x^2 +1 = 5
x^2 = 4
x = 2
x = -2